A fabrication company engaged in production of a motor part has a production capacity of 700, 000


pieces per year. But, it is just operating at 62% of its full capacity due to unavailability to finance the


importation of their materials. The company has an annual income of P 430, 000.00, annual fixed cost


are P 190, 000.00 and variable costs are P 0.348 per unit. How many productions of parts must be


produced for break-even point?

1
Expert's answer
2022-01-18T13:02:08-0500

The price is:

"P = (TP + TC)\/Q = (430,000 + 190,000 + 0.348\u00d70.62\u00d7700,000)\/(0.62\u00d7700,000) = 1.777."

The break-even point is:

"Q = 190,000\/(1.777\u00d7(1 - 0.348)) = 163,990.4" units.